### v1

# ET-Former: Efficient Triplane Deformable Attention for 3D Semantic Scene Completion From Monocular Camera

Abstract
We introduce ET-Former, a novel end-to-end algorithm for semantic scene completion using a single monocular camera. Our approach generates a semantic occupancy map from single RGB observation while simultaneously providing uncertainty estimates for semantic predictions. By designing a triplane-based deformable attention mechanism, our approach improves geometric understanding of the scene than other SOTA approaches and reduces noise in semantic predictions. Additionally, through the use of a Conditional Variational AutoEncoder (CVAE), we estimate the uncertainties of these predictions. The generated semantic and uncertainty maps will aid in the formulation of navigation strategies that facilitate safe and permissible decision-making in the future. Evaluated on the Semantic-KITTI dataset, ET-Former achieves the highest IoU and mIoU, surpassing other methods by at least $15.16\%$ in IoU and $18.12\%$ in mIoU, while reducing GPU memory usage of existing methods by $25\%-50.5\%$.
I Introduction
In robotics and autonomous driving, 3D scene understanding supports both navigation and interaction with the environment [saxena2008make3d, zhang2021holistic]. Semantic Scene Completion (SSC), also referred to as Semantic Occupancy Prediction, jointly predicts the semantic and geometric properties of an entire scene using a semantic occupancy map, including occluded regions and areas beyond the camera’s field of view (FOV) [li2023voxformer, zheng2024monoocc, cao2022monoscene]. While many approaches leverage depth cameras [choe2021volumefusion, dai2020sg] or LiDAR sensors [li2023lode, xia2023scpnet] for 3D environment perception, these sensors are often more expensive and less compact than mono-RGB cameras [cao2022monoscene]. As a result, mono-RGB cameras have gained increasing attention for 3D scene understanding in both autonomous driving and robotic navigation tasks [davison2007monoslam, li2023voxformer, zheng2024monoocc]. However, SSC using mono-cameras poses several challenges, such as accurately estimating semantics in the 3D space, predicting occluded areas outside the FOV, and get the robustness of the estimation.
Accurate estimation of pixel semantics and their real-world positions is difficult due to the inherent complexity of two key tasks [cao2022monoscene, zheng2024monoocc]: 1. semantic estimation that predicts semantic labels of the observed ares; and 2. precise projection from the image plane to the real-world environment that converts the 2D semantic areas into accurate 3D locations. From a single view, the predicted regions are strongly influenced by the pixel rays from the camera to the image plane [cao2022monoscene], often resulting in skewed geometric shapes. In such cases, multi-view perception [li2024sgs, roldao20223d] or advanced feature processing [huang2023tri, zheng2024monoocc] can help correct these errors by incorporating the geometric structure of objects from different viewpoints.
However, estimating only the visible semantic voxels is insufficient to represent a complete 3D environment, as parts of the scene may be occluded or outside the camera’s FOV [huang2023tri, cao2022monoscene]. Estimating these occluded and out-of-FOV areas requires understanding the geometric structure and semantic relationships of nearby voxels [li2023voxformer, zheng2024monoocc]. Moreover, quantifying the uncertainty in these predictions is also crucial, as it provides critical guidance for navigation strategies of safety-critical applications, like autonomous driving and rescue missions.
Figure 1: Our ET-Former predicts the semantic occupancy map from a monocular camera image to complete the scene and estimate the uncertainties of the occupancy map.
Main Contributions: In this work, we address the challenge of accurate semantic estimation by proposing a triplane-based deformable attention model that uses deformable attention mechanism [zhu2020deformable, xia2022vision] to process voxel and RGB features in three orthogonal views. Our approach provides a better geometric understanding from three views than only mono-cam frame and our invented deformable attention also reduces the computational cost of feature processing. To predict occluded and out-of-view regions, we employ a Conditional Variational AutoEncoder (CVAE) to generate semantic voxel predictions and also estimate the uncertainties of the predictions. Our approach takes advantage of the two stages idea from VoxFormer [li2023voxformer], where the first stage estimates occupancy from the mono-camera, providing guidance to the second stage for completing the 3D semantic occupancy map. Moreover, our triplane-based feature processing mechanism improves upon the deformable attention model in VoxFormer by smoothing estimation noise generated by sparse query voxels, and our CVAE generator performs more accurate prediction than other approaches. Our main contributions are as follows:
  1. 1.
    We introduce a triplane deformable model that converts 2D image features into a 3D semantic occupancy map by projecting 3D voxel queries onto three orthogonal planes, where self- / cross-deformable attentions are applied. This model significantly reduces GPU memory usage of existing methods by $25\%-50.5\%$ while refining visual and geometric information from multiple views. It achieves the highest IoU and mIoU, surpassing other methods by at least $15.16\%$ in IoU and $18.12\%$ in mIoU on the Semantic KITTI dataset.
  2. 2.
    We designed an efficient cross deformable attention model for multiple source feature processing, such as 2D image features with 3D voxel features. The deformable attention model reduces the computational cost by using flexible number of reference points in the deformable attention mechanism.
  3. 3.
    We adapt the formulation of CVAE, treating voxel queries as latent variables and reparameterizing them as Gaussian distributions to quantify uncertainty in the predicted occupancy map from stage 1, producing an uncertainty map. The integration of uncertainty and semantic maps will enable design of both safe and permissible navigation strategies in the future.
  4. 4.
    We apply the Triplane deformable model to stage 1 by fully leveraging image features to improve occupancy prediction accuracy.
II Related Works
Camera-based Scene Understanding: 3D scene understanding using cameras requires a comprehensive understanding of both geometric and color information [han2019image, li2022bevformer, huang2021bevdet], such as 3D scene reconstruction [popov2020corenet, li2023voxformer], 3D object detection [huang2023tri, huang2021bevdet], and 3D depth estimation [zhang2022dino, lin2017feature]. For tasks like detection and depth estimation, pixels and points are widely utilized to represent 3D objects [huang2021bevdet, huang2023tri, zhang2022dino, lin2017feature]. In scene reconstruction and completion tasks, early methods applied Truncated Signed Distance Functions (TSDFs) to process features [dai2020sg, park2019deepsdf]. Then voxel-based approaches gained more popularity due to their sparsity and memory efficiency [cao2022monoscene, li2023voxformer]. However, converting RGB features to 3D voxel features remains challenging because it requires both color and geometric understanding of the environment. To address this, Bird’s-Eye-View (BEV) scene representations are commonly used [li2022bevformer, zhang2022beverse]. The Lift-Splat-Shoot (LSS) method [philion2020lift] projects image features into 3D space using pixel-wise depth to generate BEV features. BEV-based methods like BevFormer [li2022bevformer] and PolarFormer [jiang2023polarformer] leverage attention mechanisms to learn the correlation between image features and 3D voxel features. These methods explicitly construct BEV features, while other works implicitly process BEV features within the network [reiser2021kilonerf, chen2021learning, huang2023tri], demonstrating higher computational efficiency and the ability to encode arbitrary scene resolutions. Recently, triplane features have been introduced to extend BEV representations [huang2023tri, shue20233d], projecting point features into three orthogonal planes instead of a single BEV plane. This approach has shown improved accuracy in 3D object detection [huang2023tri, huang2024gaussianformer]. However, these triplane-based approaches process do not process the voxels in the entire 3D scenarios. In our approach, we designed an efficient deformable-attention model to address the issue.
Semantic Scene Completion: Semantic Scene Completion (SSC) involves estimating both the geometry and semantic labels of a 3D scene, a task first introduced by SSCNet [li2023sscnet]. Early methods [chen20203d, zhou2022cross, song2017semantic, reading2021categorical] utilized geometric information to correlate 2D image features with 3D space. MonoScene [cao2022monoscene] and LMSCNet [roldao2020lmscnet] with monocular cameras proposed end-to-end solutions that convert RGB images into 3D semantic occupancy maps through convolutional neural networks (CNNs) or U-Net architectures. TPVFormer [huang2023tri] applied triplane features for pixel processing with multiple cameras but struggled with filling in missing areas effectively and heavy memory usages. More recent approaches like OccFormer [zhang2023occformer] and VoxFormer [li2023voxformer] have utilized transformers to enhance feature processing with monocular cameras. VoxFormer decouples occupancy and semantic predictions into two stages to improve accuracy in both semantic estimation and occlusion prediction. Building on VoxFormer, MonoOcc [zheng2024monoocc] and Symphonize [jiang2024symphonize] introduced more complex structures for both stages, further enhancing occupancy prediction and semantic estimation, though at the expense of higher memory usage. However, these methods do not estimate the certainty of semantic predictions, limiting their effectiveness in downstream tasks. To address this, we design a triplane-based deformable attention model to improve the estimation accuracy and incorporate Conditional Variational AutoEncoder (CVAE) method to predict the uncertainty of semantic estimations.
Figure 2: Architecture: We present a two-stage pipeline for processing mono-cam images and generate both a semantic occupancy map and its corresponding uncertainty map. In the second stage, we introduce an efficient triplane-based deformable attention mechanism, which reduces high-dimensional 3D feature processing to 2D computations. To estimate the uncertainty in the semantic map, we incorporate a CVAE method. In the first stage, we also apply the triplane-based deformable attention model to generate the occupancy queries, as shown in Figure 4.
III Our Approach
In this section, we introduce the problem definition, key innovations including the efficient triplane-based deformable 3D attention mechanism, and the Conditional Variational Autoencoder (CVAE) formulation.
III-A Problem Definition
We formulate the Semantic Scene Completion (SSC) task as a generative problem. Given a monocular image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, where $(H,W)$ represents the image size, our model aims to generate a semantic occupancy map $\mathbf{M}\in\mathbb{R}^{L\times V\times D\times N}$, where $(L,V,D)$ represents the dimensions of the 3D occupancy map and $N$ is the number of semantic classes. The model follows a CVAE formulation:
[…]
where $\mathbf{m}_{s}\in\mathbf{M}$ is the generated semantic occupancy map and $\mathbf{c}$ indicates the condition, which contains image features in our approach. The embedding vector $\mathbf{z}$ is encoded by the encoder $p_{\theta}()$, given the image instance $\mathbf{i}\in\mathbf{I}$. As illustrated in Figure 2, the model consists of a two-stage architecture for the SSC task. In the first stage, the occupancy map (queries) $\mathbf{m}_{o}\in\mathbb{R}^{L\times V\times D\times 2}$ is estimated. It serves as the inference (occupancy) for semantic prediction. In the second stage, the semantic occupancy map is generated using the CVAE formulation, with the inferred voxels from stage one as input and conditioned on the RGB image.
Our encoder in the second stage employs ResNet-50 [he2016deep] and FPN [lin2017feature], which have demonstrated high potential in image feature extraction for vision tasks [li2023voxformer, zheng2024monoocc, zhang2022dino], to process the RGB image into a lower-dimensional vector $\mathbf{f}_{r}\in\mathbb{R}^{w,h,d}$ as image features, where $w<W$, $h<H$, and $d$ denotes the feature dimension. Similar to VoxFormer [li2023voxformer], DETR [zhu2020deformable] is used to enhance the occupancy queries $\mathbf{m}_{o}$ by incorporating the image features $\mathbf{f}_{r}$ and outputs an enhanced query features $\mathbf{f}_{o}$ as follows:
$\mathbf{f}_{o}=DETR(\mathbf{z}_{q},\mathbf{p}_{o},\mathbf{f}_{r})=$
$\sum_{m=1}^{M}\mathbf{W}_{m}[\sum_{k=1}^{K}A_{mqk}$ $\mathbf{W}^{\prime}_{m}\mathbf{f}_{r}(\phi_{s}(\hat{\mathbf{p}}_{o})+\phi_{o}(\mathbf{z}_{q}))],$ (2)
where $M$ and $K$ indicate the head number and offset number. $\mathbf{p}_{o}$ represents the 3D position of each query voxel. $\mathbf{z}_{q}=\mathbf{f}_{e}(\mathbf{p}_{o})$ is the positional embedding of the query positions. $\phi_{s}$ scales the normalized position $\hat{\mathbf{p}}_{o}$ into the image feature size $h\times w$, and $\phi_{o}$ calculates the positional offset with $\mathbf{z}_{q}$. $\mathbf{W}_{m}$ and $\mathbf{W}^{\prime}_{m}$ are linear weights. $A_{mqk}=\phi_{a}(\mathbf{z}_{q})$ is also computed from $\mathbf{z}_{q}$, where $\phi_{a}()$ represents Linear layers.
III-B Conditional Variational Autoencoder Formulation
Our approach utilizes a Conditional Variational Autoencoder (CVAE) to generate semantic occupancy maps, leveraging the unique properties of the CVAE method [bao2017cvae, harvey2021conditional, sohn2015learning]: 1. CVAEs enable conditional generation, allowing for the generation of voxels that adhere to specific conditional information. We employ this capability to generate semantic maps that conform to the RGB information from monocular observations. 2. CVAEs learn a latent space conditioned on specific attributes, which is typically smooth and continuous. This property ensures that small changes in the latent space lead to meaningful variations in the generated voxels, enhancing the expressiveness of our semantic maps. 3. CVAEs inherently model uncertainty in the generated data, accommodating cases with multiple valid outputs for the same condition, unlike deterministic models. This probabilistic nature allows our approach to handle uncertainty in monocular semantic occupancy mapping more effectively.
Given the query features $\mathbf{f}_{o}$, in the CVAE format, we need to convert the feature embedding to the embedding of Gaussian distributions, which are shown in Equation 3
$\mu=\mathbf{l}_{m}(\mathbf{f}_{o}),\;\;\;\log\text{var}=\mathbf{l}_{v}(\mathbf{f}_{o}),$ (3)
where $\mathbf{l}_{m}$ and $\mathbf{l}_{v}$ are linear layers. The variance of the Gaussian distribution represents the uncertainty in the estimation $\mathbf{m}_{u}$, as depicted in Figure 2.
Figure 3: The Second Stage: We propose triplan-based efficient self-deformable attention (ESDA) and efficient cross-deformable attention (ECDA) methods in this stage to generate 3D features. The input for this stage consists of occupancy queries and image features. The query features are projected onto three orthogonal planes based on positional embedding weights. The three planes are filled by the ESDA method and then enhanced with image features, $\mathbf{f}_{r}$, by the ECDA method. Finally, the plane features are concatenated and upsampled to generate the semantic occupancy map.
Figure 4: The First Stage: We apply the novel triplane-based deformable attention method in this stage. The estimated depth pixels are projected to the 3D space as depth queries to the triplane-based deformable model. The output of the first stage is an occupancy map, the occupied voxels in which are then converted to occupancy queries.
III-C Triplane-based 3D Deformable Attention
The triplane-based representation [huang2023tri, shue20233d] encodes 3D features by leveraging three orthogonal 2D feature planes ($XY$, $XZ$, and $YZ$). This approach significantly reduces both memory and computational requirements compared to directly encoding features in dense 3D grids. In comparison to bird’s-eye-view representations, triplane encoding offers a more sufficient representation of the real-world environment. The orthogonality of the three planes provide a robust embedding of spatial features within the 3D space, making it an effective and efficient representation for complex 3D scenes. The deformable attention mechanism [zhu2020deformable, xia2022vision] is efficient in processing high-resolution features with spatial awareness of image pixels. However, current methods [zhu2020deformable, li2023voxformer] still rely on large sets of query and reference points, resulting in high computational costs for 3D feature processing. To address this, we propose an efficient triplane-based deformable attention method that enhances voxel features with image features, improving the efficiency and performance of cross-source feature extraction.
We propose an efficient deformable triplane decoder in stage two. Instead of querying the entire 3D voxel space, we reduce the dimensionality from 3D to 2D by aggregating voxel features onto three orthogonal planes. We then propose efficient self- and cross-deformable attention mechanisms to process the features of these planes. As shown in Figure 3, the query features are aggregated by the positional embedding weights:
$\mathbf{f}_{p}[x,y]=\sum_{z=0}^{Z}\mathbf{p}_{e}(\mathbf{p}_{o})\circ\mathbf{f}_{o},$ (4)
where $x$ and $y$ are the pixel indices in a plane, which also correlate to the $x,y$ in the 3D space, and $z$ represents the third axis in the 3D space. The same aggregation process also applies to other two planes. The notation $\circ$ in Eqn (4) represents element-wise product.
Since queries projected from 3D space to 2D planes only occupy part of the 2D space, there are empty pixels in the 2D space. Therefore, to have a full representation of the 3D space, we must fill in the missing pixel features. We employ the Deformable Transformer [xia2022vision] to complete the missing pixel features, using plane-wide pixel features as queries and sampling reference points at a stride of four pixels. The Efficient Self-Deformable Attention (ESDA) mechanism completes the three planes, as $\mathbf{f}_{\text{self}}=\text{ESDA}(\mathbf{f}_{p})$.
As the decoder follows the CVAE format, we leverage visual features to enhance the three orthogonal planes, where the condition is the image feature $\mathbf{f}_{r}$. We introduce a new Efficient 3D Cross Deformable Attention (ECDA) model to enhance the plane features $\mathbf{f}_{\text{self}}$ with the image feature $\mathbf{f}_{r}$. Since the plane features are converted from 3D to 2D, we retain 3D geometric representation by normally sampling reference points $\mathbf{p}_{r}\in\mathbb{R}^{l,v,d}$ from the 3D voxel grid, where $l\ll L$, $v\ll V$, and $d\ll D$. These points are projected to the image feature plane $\mathbb{R}^{w\times h}$, so we have $N_{r}=l\times v\times d$ reference points $\mathbf{p}_{i}\in\mathbb{R}^{l\times v\times d\times 2}$ in the image feature plane. The filled-in triplane features, $\mathbf{f}_{\text{self}}$, serve as queries in ECDA. Because the reference points are normally sampled from the 3D space, the $x,y$ coordinates are correlated to the triplane $x,y$ positions. We can use convolutional layers directly to calculate the offsets of the reference points, as $\Delta\mathbf{f}_{r}=\text{Conv}(\mathbf{f}_{\text{self}})$, where $\Delta\mathbf{f}_{r}\in\mathbb{R}^{l,v,c}$ and $c$ is the feature dimension. We separate the feature dimension into $d$ vectors to represent the offset of $d$ points in the $z$ axis. Finally, we have $\Delta\mathbf{p}_{i}=\mathbf{l}_{r}(\text{rearrange}(\Delta\mathbf{f}_{r}))$. Then a multi-head attention model is used to process the sampled image features with the plane features as in Equation 5
$\mathbf{f}_{\text{cross}}=\text{MultiHead}(\mathbf{f}_{\text{self}},\mathbf{f}_{r}[\mathbf{p}_{i}+\Delta\mathbf{p}_{i}]),$ (5)
where $\mathbf{f}_{\text{self}}$ is the query and $\mathbf{f}_{r}[\mathbf{p}_{i}+\Delta\mathbf{p}_{i}]$ serves as key and value of the multi-head attention model. The $\mathbf{f}_{\text{cross}}$ from three planes would be concatenated together and processed by several linear layers and upsampled to the semantic map, $\mathbf{m}_{s}=\mathbf{l}_{s}(\text{UpSample}(\mathbf{l}_{c}([\mathbf{f}_{\text{cross}}^{l},\mathbf{f}_{\text{cross}}^{h},\mathbf{f}_{\text{cross}}^{d}])))$, where $\mathbf{l}_{s}$ and $\mathbf{l}_{c}$ are all linear layers. $\mathbf{f}_{\text{cross}}^{l}$, $\mathbf{f}_{\text{cross}}^{h}$ and $\mathbf{f}_{\text{cross}}^{d}$ are features of the 2D planes orthogonal to the $l,h,d$ axes.
Methods IoU mIoU
road
sidewalk
parking
other-ground
building
car
truck
bicycle
motorcycle
other-vehicle
vegetation
trunk
terrain
person
bicyclist
motorcyclist
fence
pole
traffic-sign
MonoScene [cao2022monoscene] 34.16 11.08 54.70 27.10 24.80 5.70 14.40 18.80 3.30 0.50 0.70 4.40 14.90 2.40 19.50 1.00 1.40 0.40 11.10 3.30 2.10
TPVFormer [huang2023tri] 34.25 11.26 55.10 27.20 27.40 6.50 14.80 19.20 3.70 1.00 0.50 2.30 13.90 2.60 20.40 1.10 2.40 0.30 11.00 2.90 1.50
VoxFormer [li2023voxformer] 42.95 12.20 53.90 25.30 21.10 5.60 19.80 20.80 3.50 1.00 0.70 3.70 22.40 7.50 21.30 1.40 2.60 0.20 11.10 5.10 4.90
OccFormer [zhang2023occformer] 34.53 12.32 55.90 30.30 31.50 6.50 15.70 21.60 1.20 1.50 1.70 3.20 16.80 3.90 21.30 2.20 1.10 0.20 11.90 3.80 3.70
MonoOcc [zheng2024monoocc] - 13.80 55.20 27.80 25.10 9.70 21.40 23.20 5.20 2.20 1.50 5.40 24.00 8.70 23.00 1.70 2.00 0.20 13.40 5.80 6.40
Symphonize [jiang2024symphonize] 41.44 13.44 55.78 26.77 14.57 0.19 18.76 27.23 15.99 1.44 2.28 9.52 24.50 4.32 28.49 3.19 8.09 0.00 6.18 8.99 5.39
OctreeOcc [lu2023octreeocc] 44.71 13.12 55.13 26.74 18.68 0.65 18.69 28.07 16.43 0.64 0.71 6.03 25.26 4.89 32.47 2.25 2.57 0.00 4.01 3.72 2.36
ET-Former (Ours) 51.49 16.30 57.64 25.80 16.68 0.87 26.74 36.16 12.95 0.69 0.32 8.41 33.95 11.58 37.01 1.33 2.58 0.32 9.52 19.60 6.90
TABLE I: Semantic Scene Completion: Our ET-Former outperforms other approaches, showing at least $15.16\%$ and $18.12\%$ improvements in IoU and mIoU. For specific semantic categories, such as road, building, car, vegetation, trunk, terrain, pole, and traffic-sign, our model provides the best accuracy.
RGB Ground Truth ET-Former VoxFormer [li2023voxformer] MonoScene [cao2022monoscene]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Semantic Qualitative Results: The blue rectangles indicate missing areas where VoxFormer and MonoScene do not have but Ground Truth contains. The red rectangles marked the areas with wrong semantic estimations.
RGB Ground Truth ET-Former Uncertainty
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Uncertainty Qualitative Results: In column 3, blue voxels show low uncertainty areas, while red voxels, highlighted by yellow rectangles, indicate high uncertainty. These uncertain areas are typically outside the camera’s FOV, or obscured by vegetation (row 1) or buildings (rows 2-3).
III-D Occupancy Prediction
In the first stage, the occupancy map $\mathbf{m}_{o}$ is estimated, which identifies occupied voxels in 3D space. As shown in Figure 4, we estimate depth from the RGB image, and then the depth pixels are projected into 3D space and voxelized into an occupancy map. The initial map is often inaccurate due to raw depth estimation. Following a similar approach to stage two, we use ResNet-50 and FPN to process image features, enhancing the occupancy map with the triplane-based deformable attention model described in Section III-C. In the triplane model, we use positional embeddings as the query features and aggregate them into three orthogonal planes. The output of the triplane model is an occupancy map $\mathbf{m}_{o}=\mathbf{l}_{o}([\mathbf{f}_{\text{cross}}^{l},\mathbf{f}_{\text{cross}}^{h},\mathbf{f}_{\text{cross}}^{d}])$, where $\mathbf{l}_{o}$ is a linear layer.
III-E Training Strategy
The two stages are trained separately, using a cosine scheduler. Given that the ground truth occupancy map $\mathbf{m}_{o}^{gt}$ is derived from a sequence of LIDAR data, which may include occlusions, we mask out missing and unlabeled voxels during training. Stage one uses a cross-entropy loss $\mathcal{L}_{ce}(\mathbf{m}_{o}^{gt},\mathbf{m}_{o})$ for occupancy prediction.
The second stage is trained in the formulation of CVAE, so we have the CVAE lower bound loss function:
$\mathcal{L}_{C}=-\mathcal{D}_{KL}(p_{\theta}(\mathbf{z}|\mathbf{m}_{o},\mathbf{f}_{r})\parallel p(\mathbf{z}))+\sum\log p(\mathbf{m}_{s}|\mathbf{f}_{r},\mathbf{z}),$
where $p()$ represents the probability. Besides $\mathcal{L}_{C}$, reconstruction losses including class-weighted cross-entropy loss $\mathcal{L}_{ce}$, and losses $\mathcal{L}_{scal}^{sem}$ and $\mathcal{L}_{scal}^{geo}$ from [cao2022monoscene] are used.
IV Experiments
Dataset and Evaluation Metrics: In this work, various experiments were conducted to evaluate the proposed 3D uncertainty semantic prediction framework using the SemanticKITTI [behley2019semantickitti] dataset with monocular cameras. SemanticKITTI provides the semantic ground truth data in the range of $[0\ \text{m},51.2\ \text{m}]\times[-25.6\ \text{m},25.6\ \text{m}]\times[-2\ \text{m},4.4\ \text{m}]$. The semantic occupancy map has the shape of $256\times 256\times 32$ voxels, while the occupancy map for stage one has the shape $128\times 128\times 16$. Each voxel in the semantic occupancy map has 20 classes of semantic labels consisting of 19 semantic classes and 1 free class. Following the regular metric for SSC tasks, we utilize the Intersection over Union (IoU) to evaluate the occupancy prediction and the scene completion quality in stages one and two. For semantic evaluation, we use the mean IoU (mIoU) of the 19 semantic classes to evaluate the performance of the semantic prediction. Furthermore, we enumerate all 19 classes to evaluate the accuracy of the methods in each semantic class estimation. The training and evaluation are done by one GPU V100.
IV-A Comparisons
Table I compares our approach, ET-Former, with state-of-the-art methods. Our model achieves the highest IoU and mIoU, surpassing other methods by at least $15.16\%$ in IoU and $18.12\%$ in mIoU. The triplane-based feature representation proves particularly effective for commonly seen, large semantic objects. We achieve the best semantic accuracy across multiple categories, such as road, building, car, vegetation, trunk, terrain, pole, and traffic-sign.
We evaluated the memory usage of ET-Former against VoxFormer [li2023voxformer] and others in training memory consumption (Table II). ET-Former reduces GPU memory usage by 3.7 GB $(25\%)$ compared to VoxFormer, and by 11.1 GB $(50.5\%)$ compared to Symphonize [jiang2024symphonize], while significantly outperforming them in IoU and mIoU. These results highlight the efficiency of the proposed method.
Methods IoU mIoU Memory
VoxFormer [li2023voxformer] 42.95 12.20 14.6 Gb
MonoScene [cao2022monoscene] 34.16 11.08 18.9 Gb
Symphonize  [jiang2024symphonize] 41.44 13.44 22.0 Gb
TPVFormer [huang2023tri] 34.25 11.26 23.0 Gb
ET-Former 51.49 16.30 10.9 Gb
TABLE II: Memory Usage: Our method reduces GPU memory usage by 3.7 GB $(25\%)$ compared to VoxFormer [li2023voxformer], and by 11.1 GB $(50.5\%)$ compared to Symphonize [jiang2024symphonize], while outperforming them in IoU and mIoU.
For the evaluation of our first-stage occupancy prediction, we compared our results with those of VoxFormer, as presented in Table IV in Section VI. Our model achieves $14\%$ and $16.3\%$ improvements in IoU and Recall, respectively.
IV-B Ablation Study
Table III presents an ablation analysis of the architectural components, using VoxFormer [li2023voxformer] as the baseline. Replacing VoxFormer’s second stage with our deterministic second stage model (where query features $\mathbf{f}_{o}$ are directly fed into the deformable triplane decoder instead of converting to Gaussian distributions as in (3)) improves IoU by 0.36 and mIoU by 0.28 compared to the baseline. Implementing our CVAE stochastic second stage further enhances performance by 0.63 IoU and 0.57 mIoU. Finally, combining our first stage model with the CVAE stochastic second stage significantly boosts performance by 7.57 IoU and 3.25 mIoU.
Methods IoU mIoU
Baseline (VoxFormer [li2023voxformer]) 42.95 12.20
+ Our deterministic stage 2 43.29 (+0.36) 12.48 (+0.28)
+ Our stochastic stage 2 43.92 (+0.63) 13.05 (+0.57)
+ Our stage 1 + stochastic stage 2 51.49 (+7.57) 16.30 (+3.25)
TABLE III: Ablation Study: our key components (deformable triplane decoder, CVAE formulation, and proposed first stage model) each contribute to significant improvements in semantic estimation accuracy.
IV-C Visualizations
Qualitative Results: Figure 5 qualitatively compares our approach with VoxFormer [li2023voxformer] and MonoScene [cao2022monoscene]. Blue rectangles highlight areas missed by other methods, while red rectangles indicate incorrect semantic estimations. VoxFormer shows considerable noise in its predictions, especially outside the camera’s FOV, and struggles with occluded areas (e.g., the driving road in the third row). MonoScene has difficulty estimating semantics in occluded areas and completing the map. In contrast, our ET-Former effectively handles occluded and out-of-FOV areas.
Uncertainty Map visualization: In Figure 6, we evaluate the uncertainties in the estimated semantic occupancy map. The yellow rectangles highlight areas with high uncertainty, which mainly arise from occlusions and regions outside the camera’s FOV. For example, in the first and third rows, areas behind vegetation and buildings exhibit significant uncertainties, respectively, with the uncertainty extending outward from these objects along the camera’s line of sight. In the second row, while the horizontal road is blocked by inaccurately estimated terrain, our method assigns high uncertainty values to these problematic areas. This uncertainty quantification helps identify regions where the occupancy estimation may be less reliable, potentially aiding in formulation of risk-aware navigation strategies.
V Conclusion, Limitations, and Future Work
ET-Former offers an efficient and robust solution for 3D semantic scene completion using a single monocular camera. It successfully achieves state of the art performance with much less memory usage, highlighting its potential for real-world applications in autonomous navigation systems. Our future work could focus on designing a more memory-efficient yet effective encoder. Additionally, we aim to design a navigation strategy that makes safe but also permissible decisions leveraging both semantic and uncertainty maps generated from the proposed algorithm.

### v2

@@
 Additionally, through the use of a Conditional Variational AutoEncoder (CVAE), we estimate the uncertainties of these predictions.
-The generated semantic and uncertainty maps will aid in the formulation of navigation strategies that facilitate safe and permissible decision-making in the future.
-Evaluated on the Semantic-KITTI dataset, ET-Former achieves the highest IoU and mIoU, surpassing other methods by at least $15.16\%$ in IoU and $18.12\%$ in mIoU, while reducing GPU memory usage of existing methods by $25\%-50.5\%$.
+The generated semantic and uncertainty maps will help formulate navigation strategies that facilitate safe and permissible decision making in the future.
+Evaluated on the Semantic-KITTI dataset, ET-Former achieves the highest Intersection over Union (IoU) and mean IoU (mIoU) scores while maintaining the lowest GPU memory usage, surpassing state-of-the-art (SOTA) methods.
+It improves the SOTA scores of IoU from $44.71$ to $51.49$ and mIoU from $15.04$ to $16.30$ on SeamnticKITTI test, with a notably low training memory consumption of $10.9$ GB, achieving at least a $25\%$ reduction compared to previous methods.
+Project page: https://github.com/jingGM/ET-Former.git.
 I Introduction
@@
 Main Contributions: In this work, we address the challenge of accurate semantic estimation by proposing a triplane-based deformable attention model that uses deformable attention mechanism [zhu2020deformable, xia2022vision] to process voxel and RGB features in three orthogonal views.
-Our approach provides a better geometric understanding from three views than only mono-cam frame and our invented deformable attention also reduces the computational cost of feature processing.
+Our approach provides a better geometric understanding from three views than only mono-cam frame and our deformable attention also reduces the computational cost of feature processing.
 To predict occluded and out-of-view regions, we employ a Conditional Variational AutoEncoder (CVAE) to generate semantic voxel predictions and also estimate the uncertainties of the predictions.
-Our approach takes advantage of the two stages idea from VoxFormer [li2023voxformer], where the first stage estimates occupancy from the mono-camera, providing guidance to the second stage for completing the 3D semantic occupancy map.
+Our approach takes advantage of the two stages idea from VoxFormer [li2023voxformer], where stage 1 estimates occupancy from the mono-camera, providing guidance to stage 2 for completing the 3D semantic occupancy map.
 Moreover, our triplane-based feature processing mechanism improves upon the deformable attention model in VoxFormer by smoothing estimation noise generated by sparse query voxels, and our CVAE generator performs more accurate prediction than other approaches.
@@
     We introduce a triplane deformable model that converts 2D image features into a 3D semantic occupancy map by projecting 3D voxel queries onto three orthogonal planes, where self- / cross-deformable attentions are applied.
-This model significantly reduces GPU memory usage of existing methods by $25\%-50.5\%$ while refining visual and geometric information from multiple views.
-It achieves the highest IoU and mIoU, surpassing other methods by at least $15.16\%$ in IoU and $18.12\%$ in mIoU on the Semantic KITTI dataset.
+This model significantly reduces GPU memory usage of existing methods by at least $25\%$ while refining visual and geometric information from multiple views.
+Evaluated on the Semantic KITTI dataset, it improves the SOTA scores of IoU from $44.71$ to $51.49$ and mIoU from $15.04$ to $16.30$, with a notably low GPU memory consumption of $10.9$ GB.
   2. 2.
@@
 TPVFormer [huang2023tri] applied triplane features for pixel processing with multiple cameras but struggled with filling in missing areas effectively and heavy memory usages.
-More recent approaches like OccFormer [zhang2023occformer] and VoxFormer [li2023voxformer] have utilized transformers to enhance feature processing with monocular cameras.
+More recent approaches like OccFormer [zhang2023occformer] and VoxFormer [li2023voxformer] have utilized transformers to process features of monocular cameras.
 VoxFormer decouples occupancy and semantic predictions into two stages to improve accuracy in both semantic estimation and occlusion prediction.
-Building on VoxFormer, MonoOcc [zheng2024monoocc] and Symphonize [jiang2024symphonize] introduced more complex structures for both stages, further enhancing occupancy prediction and semantic estimation, though at the expense of higher memory usage.
+Building on VoxFormer, MonoOcc [zheng2024monoocc] and Symphonies [jiang2024symphonize] introduced more complex structures for both stages, further enhancing occupancy prediction and semantic estimation, though at the expense of higher memory usage.
 However, these methods do not estimate the certainty of semantic predictions, limiting their effectiveness in downstream tasks.
 To address this, we design a triplane-based deformable attention model to improve the estimation accuracy and incorporate Conditional Variational AutoEncoder (CVAE) method to predict the uncertainty of semantic estimations.
-Figure 2: Architecture: We present a two-stage pipeline for processing mono-cam images and generate both a semantic occupancy map and its corresponding uncertainty map.
-In the second stage, we introduce an efficient triplane-based deformable attention mechanism, which reduces high-dimensional 3D feature processing to 2D computations.
-To estimate the uncertainty in the semantic map, we incorporate a CVAE method.
-In the first stage, we also apply the triplane-based deformable attention model to generate the occupancy queries, as shown in Figure 4.
+Figure 2: Overall Architecture: We present a two-stage pipeline for processing mono-cam images and generate both a semantic occupancy map $\mathbf{m}_{s}$ and its corresponding uncertainty map $\mathbf{m}_{u}$.
+In stage 1, we introduce a novel triplane-based deformable attention model to generate the occupancy queries $\mathbf{m}_{o}$ from the given mono-cam images, which reduces high-dimensional 3D feature processing to 2D computations.
+The detailed artchitecture of stage 1 is shown in Figure 6.
+In stage 2, we employ the efficient triplane-based deformable attention mechanism (detailed in Figure 3) to generate the semantic map, with the inferred voxels from stage 1 as input and conditioned on the RGB image.
+To estimate the uncertainty in the semantic map, we incorporate a CVAE method, and quantify the uncertainty using the variance of the CVAE latent samples.
 III Our Approach
@@
 We formulate the Semantic Scene Completion (SSC) task as a generative problem.
-Given a monocular image $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$, where $(H,W)$ represents the image size, our model aims to generate a semantic occupancy map $\mathbf{M}\in\mathbb{R}^{L\times V\times D\times N}$, where $(L,V,D)$ represents the dimensions of the 3D occupancy map and $N$ is the number of semantic classes.
+Given a monocular RGB image $\mathbf{i}\in\mathbb{R}^{H\times W\times 3}$, where $(H,W)$ represents the image size, our model aims to generate a semantic occupancy map $\mathbf{M}\in\mathbb{R}^{L\times V\times D\times N}$, where $(L,V,D)$ represents the dimensions of the 3D occupancy map and $N$ is the number of semantic classes.
 The model follows a CVAE formulation:
 […]
-where $\mathbf{m}_{s}\in\mathbf{M}$ is the generated semantic occupancy map and $\mathbf{c}$ indicates the condition, which contains image features in our approach.
-The embedding vector $\mathbf{z}$ is encoded by the encoder $p_{\theta}()$, given the image instance $\mathbf{i}\in\mathbf{I}$.
-As illustrated in Figure 2, the model consists of a two-stage architecture for the SSC task.
-In the first stage, the occupancy map (queries) $\mathbf{m}_{o}\in\mathbb{R}^{L\times V\times D\times 2}$ is estimated.
-It serves as the inference (occupancy) for semantic prediction.
-In the second stage, the semantic occupancy map is generated using the CVAE formulation, with the inferred voxels from stage one as input and conditioned on the RGB image.
-Our encoder in the second stage employs ResNet-50 [he2016deep] and FPN [lin2017feature], which have demonstrated high potential in image feature extraction for vision tasks [li2023voxformer, zheng2024monoocc, zhang2022dino], to process the RGB image into a lower-dimensional vector $\mathbf{f}_{r}\in\mathbb{R}^{w,h,d}$ as image features, where $w<W$, $h<H$, and $d$ denotes the feature dimension.
-Similar to VoxFormer [li2023voxformer], DETR [zhu2020deformable] is used to enhance the occupancy queries $\mathbf{m}_{o}$ by incorporating the image features $\mathbf{f}_{r}$ and outputs an enhanced query features $\mathbf{f}_{o}$ as follows:
-$\mathbf{f}_{o}=DETR(\mathbf{z}_{q},\mathbf{p}_{o},\mathbf{f}_{r})=$
+where $\mathbf{m}_{s}\in\mathbf{M}$ is the generated semantic occupancy map given the image instance $\mathbf{i}$.
+$p(\mathbf{m}_{s}|\mathbf{z},\mathbf{i})$ denotes the conditional distribution of generating the semantic map $\mathbf{m}_{s}$ given the latent sample $\mathbf{z}$ and image $\mathbf{i}$.
+$p(\mathbf{z})$ is the prior distribution of $\mathbf{z}$.
+$p_{\theta}(\mathbf{z}|\mathbf{i})$ represents the approximate posterior distribution of the latent representation $\mathbf{z}$, given the image $\mathbf{i}$.
+As illustrated in Figure 2, the model employs of a two-stage architecture for the SSC task.
+In stage 1, the occupancy map (queries) $\mathbf{m}_{o}\in\mathbb{R}^{L\times V\times D\times 2}$ is estimated using the given image.
+The inferred $\mathbf{m}_{o}$ serves as the input for semantic prediction of stage 2.
+In stage 2, the semantic occupancy map $\mathbf{m}_{s}$ is generated using the CVAE formulation, with the inferred voxels from stage 1 as input and conditioned on the RGB image.
+As a byproduct of the CVAE formulation, reparameterizing the latent variables as Gaussian distributions allows for quantification of uncertainty in the occupancy map $\mathbf{m}_{o}$ from stage 1, giving rise to an uncertainty map $\mathbf{m}_{u}$.
+One of our major innovations is the Deformable Triplane Decoder, which is primarily designed for stage 2 but also serves as the backbone of stage 1 to provide more visual information for occupancy prediction.
+Given this interconnected design, we will first describe the methodology of stage 2 in detail.
+Subsequently, we will elaborate on the sepecifics of stage 1 in Section III-E, highlighting how the Deformable Triplane Decoder is adapted and utilized in this initial stage.
+III-B Voxel Encoder
+Our voxel encoder in stage 2 employs ResNet-50 [he2016deep] and Feed Forward Network (FFN) [lin2017feature], which have demonstrated high potential in image feature extraction for vision tasks [li2023voxformer, zheng2024monoocc, zhang2022dino], to process the RGB image into lower-dimensional features $\mathbf{f}_{r}\in\mathbb{R}^{w,h,d}$, where $w<W$, $h<H$, and $d$ denotes the feature dimension.
+DETR [zhu2020deformable] is used to enhance occupancy queries $\mathbf{m}_{o}$ by incorporating the image features $\mathbf{f}_{r}$ and outputs occupancy features $\mathbf{f}_{o}$, which is shown in Equation 2.
+This process is referred to as voxel deformable cross attention in Figure 2.
+$\mathbf{f}_{o}=DETR(\mathbf{z}_{q},\hat{\mathbf{p}}_{o},\mathbf{f}_{r})=$
 $\sum_{m=1}^{M}\mathbf{W}_{m}[\sum_{k=1}^{K}A_{mqk}$ $\mathbf{W}^{\prime}_{m}\mathbf{f}_{r}(\phi_{s}(\hat{\mathbf{p}}_{o})+\phi_{o}(\mathbf{z}_{q}))],$ (2)
-where $M$ and $K$ indicate the head number and offset number.
-$\mathbf{p}_{o}$ represents the 3D position of each query voxel.
-$\mathbf{z}_{q}=\mathbf{f}_{e}(\mathbf{p}_{o})$ is the positional embedding of the query positions.
+where $\mathbf{p}_{o}$ represents the 3D position of each query voxel in $\mathbf{m}_{o}$ and $\hat{\mathbf{p}}_{o}$ indicats its normalized position.
+$\mathbf{z}_{q}=\mathbf{p}_{e}(\mathbf{p}_{o})$ is the positional embedding of the query positions, and $\mathbf{p}_{e}()$ is a positional embedding function.
+$M$ and $K$ indicate the head number and the offset number.
 $\phi_{s}$ scales the normalized position $\hat{\mathbf{p}}_{o}$ into the image feature size $h\times w$, and $\phi_{o}$ calculates the positional offset with $\mathbf{z}_{q}$.
+$\mathbf{f}_{r}(\phi_{s}(\hat{\mathbf{p}}_{o})+\phi_{o}(\mathbf{z}_{q}))$ is the feature at location $(\phi_{s}(\hat{\mathbf{p}}_{o})+\phi_{o}(\mathbf{z}_{q}))$.
 $\mathbf{W}_{m}$ and $\mathbf{W}^{\prime}_{m}$ are linear weights.
 $A_{mqk}=\phi_{a}(\mathbf{z}_{q})$ is also computed from $\mathbf{z}_{q}$, where $\phi_{a}()$ represents Linear layers.
-III-B Conditional Variational Autoencoder Formulation
+III-C Conditional Variational Autoencoder Formulation
 Our approach utilizes a Conditional Variational Autoencoder (CVAE) to generate semantic occupancy maps, leveraging the unique properties of the CVAE method [bao2017cvae, harvey2021conditional, sohn2015learning]: 1.
@@
 2.
-CVAEs learn a latent space conditioned on specific attributes, which is typically smooth and continuous.
+CVAEs learn a smooth and continuous latent space.
 This property ensures that small changes in the latent space lead to meaningful variations in the generated voxels, enhancing the expressiveness of our semantic maps.
@@
 Given the query features $\mathbf{f}_{o}$, in the CVAE format, we need to convert the feature embedding to the embedding of Gaussian distributions, which are shown in Equation 3
-$\mu=\mathbf{l}_{m}(\mathbf{f}_{o}),\;\;\;\log\text{var}=\mathbf{l}_{v}(\mathbf{f}_{o}),$ (3)
-where $\mathbf{l}_{m}$ and $\mathbf{l}_{v}$ are linear layers. The variance of the Gaussian distribution represents the uncertainty in the estimation $\mathbf{m}_{u}$, as depicted in Figure 2.
-Figure 3: The Second Stage: We propose triplan-based efficient self-deformable attention (ESDA) and efficient cross-deformable attention (ECDA) methods in this stage to generate 3D features.
-The input for this stage consists of occupancy queries and image features.
-The query features are projected onto three orthogonal planes based on positional embedding weights.
-The three planes are filled by the ESDA method and then enhanced with image features, $\mathbf{f}_{r}$, by the ECDA method.
+$\hat{\mathbf{f}}_{o}\sim\mathcal{N}\{\mu=\mathbf{l}_{m}(\mathbf{f}_{o}),\;\;\;\log\text{var}=\mathbf{l}_{v}(\mathbf{f}_{o})\},$ (3)
+where $\mathbf{l}_{m}$ and $\mathbf{l}_{v}$ are linear layers.
+$\mu$ and $var$ are the mean and variance of the distribution.
+The variance of the Gaussian distribution represents the uncertainty in the uncertainty map $\mathbf{m}_{u}$, as shown in Figure 2.
+Figure 3: (a) Deformable Triplane Decoder of Stage 2: We propose triplane-based efficient self-deformable attention (ESDA) and efficient cross-deformable attention (ECDA) methods in the decoder to generate 3D features.
+The input to the decoder consists of Gaussian samples $\hat{\mathbf{f}}_{o}$ and image features $\mathbf{f}_{r}$.
+The Gaussian samples are projected onto three orthogonal planes based on positional embeddings.
+The three planes features are completed by ESDA and then enhanced with image features, $\mathbf{f}_{r}$, by ECDA.
 Finally, the plane features are concatenated and upsampled to generate the semantic occupancy map.
-Figure 4: The First Stage: We apply the novel triplane-based deformable attention method in this stage.
-The estimated depth pixels are projected to the 3D space as depth queries to the triplane-based deformable model.
-The output of the first stage is an occupancy map, the occupied voxels in which are then converted to occupancy queries.
-III-C Triplane-based 3D Deformable Attention
+(b) Detailed Structure of ECDA: ECDA utilizes image features at deformed points $(\mathbf{p}_{i}+\Delta\mathbf{p}_{i})$ and completed plane feature $\mathbf{f}_{\text{self}}$ to conduct cross attention.
+III-D Triplane-based 3D Deformable Attention
 The triplane-based representation [huang2023tri, shue20233d] encodes 3D features by leveraging three orthogonal 2D feature planes ($XY$, $XZ$, and $YZ$).
@@
 However, current methods [zhu2020deformable, li2023voxformer] still rely on large sets of query and reference points, resulting in high computational costs for 3D feature processing.
-To address this, we propose an efficient triplane-based deformable attention method that enhances voxel features with image features, improving the efficiency and performance of cross-source feature extraction.
-We propose an efficient deformable triplane decoder in stage two.
+To address this, we propose an efficient triplane-based deformable attention method that enhances voxel features with image features $\mathbf{f}_{r}$, improving the efficiency and performance of cross-source feature extraction.
+We propose an efficient deformable triplane decoder in stage 2.
 Instead of querying the entire 3D voxel space, we reduce the dimensionality from 3D to 2D by aggregating voxel features onto three orthogonal planes.
 We then propose efficient self- and cross-deformable attention mechanisms to process the features of these planes.
-As shown in Figure 3, the query features are aggregated by the positional embedding weights:
-$\mathbf{f}_{p}[x,y]=\sum_{z=0}^{Z}\mathbf{p}_{e}(\mathbf{p}_{o})\circ\mathbf{f}_{o},$ (4)
-where $x$ and $y$ are the pixel indices in a plane, which also correlate to the $x,y$ in the 3D space, and $z$ represents the third axis in the 3D space.
+As shown in Figure 3, the query features $\hat{\mathbf{f}}_{o}$ are aggregated by the positional embedding weights:
+$\mathbf{f}_{p}(x,y)=\sum_{z=0}^{Z}\mathbf{p}_{e}(\mathbf{p}_{o}(x,y,z))\circ\hat{\mathbf{f}}_{o}(x,y,z),$ (4)
+where $x$ and $y$ are the cell indices in a plane, which also correlate to the $x,y$ in the 3D space, and $z$ represents the third axis in the 3D space.
+$\mathbf{p}_{e}$ is a positional embedding function.
+$\hat{\mathbf{f}}_{o}$ is a sample from the normal distribution in Equation 3.
+The notation $\circ$ in Equation 4 represents element-wise product.
+In short, $\mathbf{f}_{p}$ at location $(x,y)$ is the weighted sum of $\hat{\mathbf{f}}_{o}(x,y,z)$ across all values of $z$.
 The same aggregation process also applies to other two planes.
-The notation $\circ$ in Eqn (4) represents element-wise product.
-Since queries projected from 3D space to 2D planes only occupy part of the 2D space, there are empty pixels in the 2D space.
-Therefore, to have a full representation of the 3D space, we must fill in the missing pixel features.
-We employ the Deformable Transformer [xia2022vision] to complete the missing pixel features, using plane-wide pixel features as queries and sampling reference points at a stride of four pixels.
+Since queries projected from 3D space to 2D planes only occupy part of the 2D space, there are empty cells in the 2D space.
+Therefore, to have a full representation of the 3D space, we must fill in the missing cell features.
+As for the SemanicKITTI dataset [behley2019semantickitti], the missing cells are more than half of the plane, so the plane features are sparse.
+We employ the Deformable Transformer [xia2022vision] to complete the missing cell features, using plane-wide cell features as queries.
+The reference points are downsampled from the image feature space at a stride of four pixels to strike a balance between computation efficiency and performance.
 The Efficient Self-Deformable Attention (ESDA) mechanism completes the three planes, as $\mathbf{f}_{\text{self}}=\text{ESDA}(\mathbf{f}_{p})$.
-As the decoder follows the CVAE format, we leverage visual features to enhance the three orthogonal planes, where the condition is the image feature $\mathbf{f}_{r}$.
-We introduce a new Efficient 3D Cross Deformable Attention (ECDA) model to enhance the plane features $\mathbf{f}_{\text{self}}$ with the image feature $\mathbf{f}_{r}$.
-Since the plane features are converted from 3D to 2D, we retain 3D geometric representation by normally sampling reference points $\mathbf{p}_{r}\in\mathbb{R}^{l,v,d}$ from the 3D voxel grid, where $l\ll L$, $v\ll V$, and $d\ll D$.
-These points are projected to the image feature plane $\mathbb{R}^{w\times h}$, so we have $N_{r}=l\times v\times d$ reference points $\mathbf{p}_{i}\in\mathbb{R}^{l\times v\times d\times 2}$ in the image feature plane.
+Since the decoder follows the CVAE format, we leverage image features, $\mathbf{f}_{r}$, as a condition to provide different view information and color information, thereby enhancing the features of the three orthogonal planes.
+We introduce a new Efficient 3D Cross Deformable Attention (ECDA) model to process the plane features $\mathbf{f}_{\text{self}}$ with the image feature $\mathbf{f}_{r}$, which is illustrated in Figure 3.
+Since the plane features are converted from 3D to 2D, we retain 3D geometric representation by uniformly sampling reference points $\mathbf{p}_{r}\in\mathbb{R}^{l,v,d}$ from the 3D voxel grid, where $l\ll L$, $v\ll V$, and $d\ll D$.
+These points in the camera’s view, can be projected to the image feature plane $\mathbb{R}^{w\times h}$, so we have $N_{r}=l\times v\times d$ reference points $\mathbf{p}_{i}\in\mathbb{R}^{l\times v\times d\times 2}$ in the image feature plane.
 The filled-in triplane features, $\mathbf{f}_{\text{self}}$, serve as queries in ECDA.
-Because the reference points are normally sampled from the 3D space, the $x,y$ coordinates are correlated to the triplane $x,y$ positions.
-We can use convolutional layers directly to calculate the offsets of the reference points, as $\Delta\mathbf{f}_{r}=\text{Conv}(\mathbf{f}_{\text{self}})$, where $\Delta\mathbf{f}_{r}\in\mathbb{R}^{l,v,c}$ and $c$ is the feature dimension.
+Because the reference points are uniformly sampled from the 3D space, the $x,y$ coordinates are correlated to the triplane $x,y$ positions.
+We can use convolutional layers directly to calculate the offsets of the reference points, as $\Delta\mathbf{p}_{r}=\text{Conv}(\mathbf{f}_{\text{self}})$, where $\Delta\mathbf{p}_{r}\in\mathbb{R}^{l\times v\times c}$ and $c$ is the feature dimension.
 We separate the feature dimension into $d$ vectors to represent the offset of $d$ points in the $z$ axis.
-Finally, we have $\Delta\mathbf{p}_{i}=\mathbf{l}_{r}(\text{rearrange}(\Delta\mathbf{f}_{r}))$.
-Then a multi-head attention model is used to process the sampled image features with the plane features as in Equation 5
-$\mathbf{f}_{\text{cross}}=\text{MultiHead}(\mathbf{f}_{\text{self}},\mathbf{f}_{r}[\mathbf{p}_{i}+\Delta\mathbf{p}_{i}]),$ (5)
-where $\mathbf{f}_{\text{self}}$ is the query and $\mathbf{f}_{r}[\mathbf{p}_{i}+\Delta\mathbf{p}_{i}]$ serves as key and value of the multi-head attention model.
+Finally, we have $\Delta\mathbf{p}_{i}=\mathbf{l}_{r}(s(\Delta\mathbf{p}_{r}))$, where $s()$ represents the feature separation, and $\mathbf{l}_{r}$ is a linear layer.
+Then a multi-head cross attention model is used to process the sampled image features with the plane features as in Equation 5:
+$\mathbf{f}_{\text{cross}}=\text{CrossAttention}(\mathbf{f}_{\text{self}},\mathbf{f}_{r}(\mathbf{p}_{i}+\Delta\mathbf{p}_{i})),$ (5)
+where $\mathbf{f}_{\text{self}}$ is the query.
+$\mathbf{f}_{r}(\mathbf{p}_{i}+\Delta\mathbf{p}_{i})$, the image feature at deformed point locations $(\mathbf{p}_{i}+\Delta\mathbf{p}_{i})$, serves as key and value of the multi-head attention model.
 The $\mathbf{f}_{\text{cross}}$ from three planes would be concatenated together and processed by several linear layers and upsampled to the semantic map, $\mathbf{m}_{s}=\mathbf{l}_{s}(\text{UpSample}(\mathbf{l}_{c}([\mathbf{f}_{\text{cross}}^{l},\mathbf{f}_{\text{cross}}^{h},\mathbf{f}_{\text{cross}}^{d}])))$, where $\mathbf{l}_{s}$ and $\mathbf{l}_{c}$ are all linear layers.
@@
 MonoOcc [zheng2024monoocc] - 13.80 55.20 27.80 25.10 9.70 21.40 23.20 5.20 2.20 1.50 5.40 24.00 8.70 23.00 1.70 2.00 0.20 13.40 5.80 6.40
-Symphonize [jiang2024symphonize] 41.44 13.44 55.78 26.77 14.57 0.19 18.76 27.23 15.99 1.44 2.28 9.52 24.50 4.32 28.49 3.19 8.09 0.00 6.18 8.99 5.39
+Symphonies [jiang2024symphonize] 42.19 15.04 58.40 29.30 26.90 11.70 24.70 23.60 3.20 3.60 2.60 5.60 24.20 10.00 23.10 3.20 1.90 2.00 16.10 7.70 8.00
 OctreeOcc [lu2023octreeocc] 44.71 13.12 55.13 26.74 18.68 0.65 18.69 28.07 16.43 0.64 0.71 6.03 25.26 4.89 32.47 2.25 2.57 0.00 4.01 3.72 2.36
 ET-Former (Ours) 51.49 16.30 57.64 25.80 16.68 0.87 26.74 36.16 12.95 0.69 0.32 8.41 33.95 11.58 37.01 1.33 2.58 0.32 9.52 19.60 6.90
-TABLE I: Semantic Scene Completion: Our ET-Former outperforms other approaches, showing at least $15.16\%$ and $18.12\%$ improvements in IoU and mIoU.
-For specific semantic categories, such as road, building, car, vegetation, trunk, terrain, pole, and traffic-sign, our model provides the best accuracy.
+TABLE I: Quantitative Comparison On SemanticKITTI test [behley2019semantickitti]: Our ET-Former outperforms other camera-based approaches, improving the SOTA scores of IoU from $44.71$ to $51.49$ and mIoU from $15.04$ to $16.30$.
+For specific semantic categories, such as building, car, other-vehicle, vegetation, trunk, terrain, and pole, our model provides the best accuracy.
 RGB Ground Truth ET-Former VoxFormer [li2023voxformer] MonoScene [cao2022monoscene]
@@
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 5: Semantic Qualitative Results: The blue rectangles indicate missing areas where VoxFormer and MonoScene do not have but Ground Truth contains.
-The red rectangles marked the areas with wrong semantic estimations.
+Figure 4: Semantic Qualitative Results: The blue rectangles highlight areas that are present in the Ground Truth but are missing from both VoxFormer and MonoScene predictions.
+The red rectangles marked the areas with wrong semantic estimations from Voxformer and MonoScene.
 RGB Ground Truth ET-Former Uncertainty
@@
 Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 6: Uncertainty Qualitative Results: In column 3, blue voxels show low uncertainty areas, while red voxels, highlighted by yellow rectangles, indicate high uncertainty.
+Figure 5: Uncertainty Qualitative Results: In column 4, blue voxels show low uncertainty areas, while red voxels, highlighted by yellow rectangles, indicate high uncertainty.
 These uncertain areas are typically outside the camera’s FOV, or obscured by vegetation (row 1) or buildings (rows 2-3).
-III-D Occupancy Prediction
-In the first stage, the occupancy map $\mathbf{m}_{o}$ is estimated, which identifies occupied voxels in 3D space.
-As shown in Figure 4, we estimate depth from the RGB image, and then the depth pixels are projected into 3D space and voxelized into an occupancy map.
+III-E Occupancy Prediction
+In stage 1, the occupancy map $\mathbf{m}_{o}$ is estimated, which identifies occupied voxels in 3D space.
+As shown in Figure 6, we estimate depth from the RGB image, and then the depth pixels are projected into 3D space, voxelized into a raw occupancy map, and finally converted to raw occupancy queries through positional embedding.
 The initial map is often inaccurate due to raw depth estimation.
-Following a similar approach to stage two, we use ResNet-50 and FPN to process image features, enhancing the occupancy map with the triplane-based deformable attention model described in Section III-C.
-In the triplane model, we use positional embeddings as the query features and aggregate them into three orthogonal planes.
+Following a similar approach to stage 2, we use ResNet-50 and FPN to process image features, enhancing the occupancy map with the triplane-based deformable attention model described in Section III-D.
+In the triplane model, we use raw occupancy query as the query features and aggregate them into three orthogonal planes.
 The output of the triplane model is an occupancy map $\mathbf{m}_{o}=\mathbf{l}_{o}([\mathbf{f}_{\text{cross}}^{l},\mathbf{f}_{\text{cross}}^{h},\mathbf{f}_{\text{cross}}^{d}])$, where $\mathbf{l}_{o}$ is a linear layer.
-III-E Training Strategy
+Figure 6: Stage 1: We employ the novel triplane-based deformable attention method in this stage.
+The estimated depth pixels are projected to the 3D space as raw occupancy queries to the triplane-based deformable model.
+The output of stage 1 is an occupancy map, the occupied voxels in which are then converted to occupancy queries.
+III-F Training Strategy
 The two stages are trained separately, using a cosine scheduler.
 Given that the ground truth occupancy map $\mathbf{m}_{o}^{gt}$ is derived from a sequence of LIDAR data, which may include occlusions, we mask out missing and unlabeled voxels during training.
-Stage one uses a cross-entropy loss $\mathcal{L}_{ce}(\mathbf{m}_{o}^{gt},\mathbf{m}_{o})$ for occupancy prediction.
-The second stage is trained in the formulation of CVAE, so we have the CVAE lower bound loss function:
+Stage 1 uses a cross-entropy loss $\mathcal{L}_{ce}(\mathbf{m}_{o}^{gt},\mathbf{m}_{o})$ for occupancy prediction.
+Stage 2 is trained in the formulation of CVAE, so we have the CVAE lower bound loss function:
 $\mathcal{L}_{C}=-\mathcal{D}_{KL}(p_{\theta}(\mathbf{z}|\mathbf{m}_{o},\mathbf{f}_{r})\parallel p(\mathbf{z}))+\sum\log p(\mathbf{m}_{s}|\mathbf{f}_{r},\mathbf{z}),$
-where $p()$ represents the probability.
-Besides $\mathcal{L}_{C}$, reconstruction losses including class-weighted cross-entropy loss $\mathcal{L}_{ce}$, and losses $\mathcal{L}_{scal}^{sem}$ and $\mathcal{L}_{scal}^{geo}$ from [cao2022monoscene] are used.
+where $p_{\theta}(\mathbf{z}|\mathbf{m}_{o},\mathbf{f}_{r})$ represents the approximate posterior distribution parameterized by the Voxel Encoder, and the KL divergence trains the distribution close to, $p(\mathbf{z})$, the Gaussian distribution of $\mathbf{z}$.
+Note that the posterior distribution $p_{\theta}(\mathbf{z}|\mathbf{m}_{o},\mathbf{f}_{r})$ was used to sample $\hat{\mathbf{f}}_{o}$ during training as Equation 3.
+$p(\mathbf{m}_{s}|\mathbf{f}_{r},\mathbf{z})$ represents the conditional distribution of generating the semantic map $\mathbf{m}_{s}$ from the latent sample $\mathbf{z}$ and image feature $\mathbf{f}_{r}$.
+Besides KL divergence, the reconstruction loss includes class-weighted cross-entropy loss $\mathcal{L}_{ce}$, and losses $\mathcal{L}_{scal}^{sem}$ and $\mathcal{L}_{scal}^{geo}$ from [cao2022monoscene].
 IV Experiments
@@
 SemanticKITTI provides the semantic ground truth data in the range of $[0\ \text{m},51.2\ \text{m}]\times[-25.6\ \text{m},25.6\ \text{m}]\times[-2\ \text{m},4.4\ \text{m}]$.
-The semantic occupancy map has the shape of $256\times 256\times 32$ voxels, while the occupancy map for stage one has the shape $128\times 128\times 16$.
+The semantic occupancy map has the shape of $256\times 256\times 32$ voxels, while the occupancy map for stage 1 has the shape $128\times 128\times 16$.
 Each voxel in the semantic occupancy map has 20 classes of semantic labels consisting of 19 semantic classes and 1 free class.
@@
 Table I compares our approach, ET-Former, with state-of-the-art methods.
-Our model achieves the highest IoU and mIoU, surpassing other methods by at least $15.16\%$ in IoU and $18.12\%$ in mIoU.
+Our model achieves the highest IoU and mIoU, surpassing other methods by improving the SOTA scores of IoU by $15.16\%$ (from $44.71$ to $51.49$) and mIoU by $8.38\%$ (from $15.04$ to $16.30$).
 The triplane-based feature representation proves particularly effective for commonly seen, large semantic objects.
-We achieve the best semantic accuracy across multiple categories, such as road, building, car, vegetation, trunk, terrain, pole, and traffic-sign.
-We evaluated the memory usage of ET-Former against VoxFormer [li2023voxformer] and others in training memory consumption (Table II).
-ET-Former reduces GPU memory usage by 3.7 GB $(25\%)$ compared to VoxFormer, and by 11.1 GB $(50.5\%)$ compared to Symphonize [jiang2024symphonize], while significantly outperforming them in IoU and mIoU.
+We achieve the best semantic accuracy across multiple categories, such as building, car, other-vehicle, vegetation, trunk, terrain, and pole.
+We evaluated the training memory usage of ET-Former against other approaches (Table II).
+ET-Former reduces GPU memory usage by 3.7 GB $(25\%)$ compared to VoxFormer, and by 11.1 GB $(50.5\%)$ compared to Symphonies [jiang2024symphonize], while significantly outperforming them in IoU and mIoU.
 These results highlight the efficiency of the proposed method.
@@
 MonoScene [cao2022monoscene] 34.16 11.08 18.9 Gb
-Symphonize  [jiang2024symphonize] 41.44 13.44 22.0 Gb
+Symphonies  [jiang2024symphonize] 42.19 15.04 22.0 Gb
 TPVFormer [huang2023tri] 34.25 11.26 23.0 Gb
 ET-Former 51.49 16.30 10.9 Gb
-TABLE II: Memory Usage: Our method reduces GPU memory usage by 3.7 GB $(25\%)$ compared to VoxFormer [li2023voxformer], and by 11.1 GB $(50.5\%)$ compared to Symphonize [jiang2024symphonize], while outperforming them in IoU and mIoU.
-For the evaluation of our first-stage occupancy prediction, we compared our results with those of VoxFormer, as presented in Table IV in Section VI.
-Our model achieves $14\%$ and $16.3\%$ improvements in IoU and Recall, respectively.
+TABLE II: Training Memory Usage On SemanticKITTI: Our method reduces GPU memory usage by 3.7 GB $(25\%)$ compared to VoxFormer [li2023voxformer], and by 11.1 GB $(50.5\%)$ compared to Symphonies [jiang2024symphonize], while outperforming them in IoU and mIoU.
+To evaluate our stage 1 occupancy prediction model proposed in Section III-E, we conducted a comparative analysis with VoxFormer [li2023voxformer].
+To ensure a fair comparison, we utilize the same depth estimation method, MobileStereoNet [shamsafar2022mobilestereonet], as employed by the stage 1 of VoxFormer.
+Subsequently, as depicted in Figure 6, we apply our triplane-based method to generate the occupancy map and queries.
+The quality of the occupancy map predictions is compared between our stage 1 approach and VoxFormer’s, as shown in Table III.
+Our approach shows substantial improvements, with a $14.2\%$ increase in IoU and a $16.3\%$ increase in Recall, resulting in a more accurate occupancy map and more efficient occupancy queries.
+Figure 7 shows the qualitative comparison between our stage 1 approach and VoxFormer’s.
+We observe that our approach, ET-Former, can estimate details more accurately and generate more complete occupancy map than VoxFormer.
+The improvement in stage 1 ultimately contributes to the performance improvement of the subsequent semantic predictions, as shown in Table IV.
+Methods IoU Recall
+VoxFormer’s stage 1 52.03 61.52
+Our stage 1 59.41 71.56
+TABLE III: Results of Stage 1 On SemanticKITTI test [behley2019semantickitti]: Compared with the VoxFormer’s stage 1 model, our stage 1 model achieves $14.2\%$ and $16.3\%$ improvements in IoU and Recall values, respectively.
+RGB
+Refer to caption Refer to caption
+VoxFormer [li2023voxformer]
+Refer to caption Refer to caption
+ET-Former
+Refer to caption Refer to caption
+Figure 7: Stage 1 Restuls: The major differences between the stage 1 results of ET-Former and VoxFormer are marked by red rectangles.
+In the first column, ET-Former demonstrates superior performance in detecting trees and cars, providing more detailed and accurate representations.
+The second column illustrates ET-Former’s ability to generate a more comprehensive map, including the detection of distant roads that VoxFormer has missed.
 IV-B Ablation Study
-Table III presents an ablation analysis of the architectural components, using VoxFormer [li2023voxformer] as the baseline.
-Replacing VoxFormer’s second stage with our deterministic second stage model (where query features $\mathbf{f}_{o}$ are directly fed into the deformable triplane decoder instead of converting to Gaussian distributions as in (3)) improves IoU by 0.36 and mIoU by 0.28 compared to the baseline.
-Implementing our CVAE stochastic second stage further enhances performance by 0.63 IoU and 0.57 mIoU.
-Finally, combining our first stage model with the CVAE stochastic second stage significantly boosts performance by 7.57 IoU and 3.25 mIoU.
+Table IV presents an ablation analysis of the architectural components, using VoxFormer [li2023voxformer] as the baseline.
+Replacing VoxFormer’s stage 2 with our deterministic stage 2 model (where query features $\mathbf{f}_{o}$ are directly fed into the deformable triplane decoder instead of converting to Gaussian samples $\hat{\mathbf{f}}_{o}$ as in Equation 3) improves IoU by 0.36 and mIoU by 0.28 compared to the baseline.
+Implementing our CVAE stochastic stage 2 further improves performance by 0.63 IoU and 0.57 mIoU.
+Finally, combining our stage 1 model with the CVAE stochastic stage 2 significantly boosts performance by 7.57 IoU and 3.25 mIoU.
 Methods IoU mIoU
@@
 + Our stage 1 + stochastic stage 2 51.49 (+7.57) 16.30 (+3.25)
-TABLE III: Ablation Study: our key components (deformable triplane decoder, CVAE formulation, and proposed first stage model) each contribute to significant improvements in semantic estimation accuracy.
+TABLE IV: Ablation Study On SemanticKITTI test [behley2019semantickitti]: our key components (deformable triplane decoder, CVAE formulation, and proposed stage 1 model) each contribute to significant improvements in semantic estimation accuracy.
 IV-C Visualizations
 Qualitative Results: Figure 4 qualitatively compares our approach with VoxFormer [li2023voxformer] and MonoScene [cao2022monoscene].
-Blue rectangles highlight areas missed by other methods, while red rectangles indicate incorrect semantic estimations.
+Blue rectangles highlight areas present in the Ground Truth but missed by VoxFormer and MonoScene.
+Red rectangles indicate regions where VoxFormer and MonoScene made incorrect semantic estimations.
 VoxFormer shows considerable noise in its predictions, especially outside the camera’s FOV, and struggles with occluded areas (e.g., the driving road in the third row).
@@
 ET-Former offers an efficient and robust solution for 3D semantic scene completion using a single monocular camera.
-It successfully achieves state of the art performance with much less memory usage, highlighting its potential for real-world applications in autonomous navigation systems.
+It successfully achieves state-of-the-art performance with much less memory usage, highlighting its potential for real-world applications in autonomous navigation systems.
 Our future work could focus on designing a more memory-efficient yet effective encoder.